the width of a rectangle is 11 inches less than its length. find the dimensions of the rectangle if the area is 80 square inches.
Answer by unlockmath(1688)

Hello,
Let's set up x as the length and x-11 as the width. Now we can set up an equation like:
x(x-11)=80 Rewritten as:
x^2-11x=80 Subtract 80 from both sides to get:
x^2-11x-80=0 Factor this to get:
(x-16)(x+5)=0 Solve for x
x=16
x=-5 OK, we know this doesn't work,
so the length is:
16 inches and the width is 5 inches. Check it out to see if it works.
There's a book I wrote that might help:
www.math-unlock.com